no. of molecules of co2 in 44mg of co2 is
1)6.023 * 10^23
2)6.023 * 10^22
3)6.023 * 10^21
4)6.023 * 10^20

Grade:12th pass

2 Answers

Arun
5 years ago
First we get and calculate the atomic weight of carbon and oxygen
Carbon = 12 and oxygen=16
CO2=12+(16×2)
=12+32
=44u
44gm of CO2 has 6.02×10^23 molecules (Avogadro's number)
Therefore, 44g=(44×6.02×10^23)/44
=6.02×10^23 molecules
 
hence in 44mg = 6.02 x 10^21 molecules
 
Darsh patel
5 years ago
The molecular mass of CO2 is 44 gm.Here given mass is 44mg .So no of moles of CO2 is (44*10^-3)gm/44 gm which gives us 10^-3 moles of CO2.Now as CO2 is a molecule itself.So no of molecules has formula (moles*Avogadro no.) Which gives us the no of molecules as[( 10^-3)*6.022*(10^23) molecules given that Avogadro no is equal to 6.022*10^23 .So the answer is 6.022*10^20 molecules going with option D.plz approve my answer.Best of luck for your preparation.